OpenStudy (anonymous):

helkp Write the equation of a line in slope-intercept form that has a slope of -1/4 and passes through the point (8, -1)

OpenStudy (anonymous):

@philo1234

OpenStudy (anonymous):

Slope intercept form is y=mx+b, where m= slope and b is the y-intercept

OpenStudy (anonymous):

\[y-(-1)=-\frac{ 1 }{ 4 }(x-8)\] point slope form now solve for y

OpenStudy (anonymous):

Use the point-slope form: y-y1=m(x-x1), where y1=-1 and x1=8, m=-1/4

OpenStudy (anonymous):

idk how to do tht?

OpenStudy (anonymous):

just plug in the values that u got and then put y by itself on one side.

OpenStudy (anonymous):

i still dnt get it?

OpenStudy (anonymous):

y+1=-1/4(x-8), then solve for 'y'

OpenStudy (anonymous):

y+1=-1/4(x-8) -1

OpenStudy (anonymous):

Solve for 'y' u know how to do that right?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

@philo1234 ????

OpenStudy (anonymous):

I would suggest u get rid of the paretheses on the right hand side by multiplying each term in the paretheses by -1/4 . then subtract 1 from both sides

OpenStudy (anonymous):

y+1=-1/4 -1 -1

OpenStudy (anonymous):

now wat @philo1234

OpenStudy (anonymous):

I know u r messing with me. I think u know what to do. But here it is anyways: y+1 = -1/4x-8(-1/4) y+1 = -1/4x+2 -1 -1 y = -1/4x+1

OpenStudy (anonymous):

umm no i dont lol im just not smart 0.o